# Structure of the mean-field yrast spectrum of a two-component Bose gas in a ring: role of interaction asymmetry

### v1

@@ Abstract
The mean-field yrast spectrum of an SU(2)-symmetric two-component Bose gas confined to a ring geometry is known to exhibit an intricate nonanalytic structure that is absent in single-component systems. In particular, due to the interplay between the species concentration and the atomic interactions, a sequence of plane-wave states can emerge as yrast states at fractional values of the angular momentum per particle. This behavior stands in sharp contrast to the single-component case, where plane-wave states occur only at integer angular momenta. In this paper, we investigate how the structure of the yrast spectrum in a two-component Bose gas is modified by interaction asymmetry. By numerically solving the coupled Gross–Pitaevskii equations for propagating soliton states, we compute the mean-field yrast spectrum and, in particular, determine the critical curves associated with the emergence of various plane-wave yrast states. We find that both the behavior of these critical curves and the mechanisms by which plane-wave yrast states arise depend sensitively on the relative strengths of the inter- and intra-component interactions. When the inter-component interaction is weaker, the plane-wave yrast states replace soliton states through a continuous evolution, as in the SU(2)-symmetric case, although the conditions for their existence become more restrictive. In contrast, when the inter-component interaction is stronger, plane-wave yrast states emerge by overtaking soliton states via branch crossings, and their stability is significantly enhanced. Our results have important implications for the existence and stability of persistent currents in asymmetric, two-component Bose gases.
@@ I Introduction
In recent decades, quantum gases have become as an important platform for investigating persistent currents for two principal reasons [Ryu07, Ramanathan11, Moulder12, Beattie13, Wright13, Cai22, Del22, Polo25]. First, experimental advances in realizing multicomponent superfluids and synthetic gauge fields have substantially broadened the class of systems that may support persistent currents [KAWAGUCHI12, Lin_2016]. Second, the yrast spectrum of such systems can be determined quantitatively by mean-field theory, revealing a rich structure and enabling persistent currents beyond the conventional single-component scenario [Smyrnakis09, Anoshkin2013, Wu2013, Abad14, Wu2015, Pecci23]. Indeed, the mean-field yrast spectrum of a SU(2)-symmetric two-component Bose gas in a ring geometry is known to exhibit an intricate structure not present in single-component systems [Smyrnakis09, Anoshkin2013, Wu2013]. Although it still obeys the general form given in Eq. (1), the yrast spectrum of this two-component system can develop a series of non-analytic points at certain fractional values of $l$, arising from the interplay between component concentration and atomic interactions. The states at these additional points, much like those at integer $l$, may also support persistent currents. This feature of the yrast spectrum was first uncovered using analytic soliton solutions of the coupled Gross–Pitaevskii equations for the SU(2)-symmetric system, and its implications for persistent currents were later confirmed experimentally [Beattie13]. Notably, the appearance of each additional non-analytic point at the fractional $l$ is accompanied by a transition in the condensate wave functions: a pair of soliton solutions is replaced by a pair of plane-wave states. Moreover, these plane-wave states carry different winding numbers, reflecting distinct quantizations of circulation in the two components.
@@
Most experimentally accessible two-component systems are not SU(2) symmetric, since the intra- and inter-component interaction strengths generally differ. In this case, the coupled Gross–Pitaevskii (GP) equations no longer admit analytic solutions, and it is therefore natural to ask whether the recently discovered features in the yrast spectrum persist under these more realistic conditions. Reference [Wu2015] attempted to address this question using a perturbative approach based on an assumption motivated by the SU(2)-symmetric case—namely, that the appearance of a nonanalytic point at a fractional value of the angular momentum $l$ is always accompanied by a continuous transformation of the corresponding soliton state into a plane-wave state at that angular momentum. However, this assumption has never been verified for asymmetric systems. Several other works [Smyrnakis2014, Sandin16, Matsushita2018, Roussou2018] have also numerically investigated the yrast spectrum of asymmetric two-component systems, but an overarching picture of the structure of the yrast spectrum has yet to emerge. In this work, we present a comprehensive study of the mean-field yrast spectrum of asymmetric two-component Bose gases confined to a ring geometry by numerically solving the coupled GP equations. We find that the assumption adopted in Ref. [Wu2015] holds only for asymmetric two-component gases in which the inter-component interaction is weaker than the intra-component interaction. In the opposite regime, this assumption breaks down: the plane-wave yrast states instead emerge from crossings between distinct branches of solutions to the GP equations, and the resulting yrast spectrum exhibits an even richer nonanalytic structure than in the SU(2)-symmetric case. In both regimes, we establish critical conditions for the emergence of plane-wave yrast states, which to a large extent determine the analytic structure of the yrast spectrum.
@@ II Statement of the problem
We consider a two-component Bose gas of $N$ atoms confined to a ring of radius $R$, where $N_{A}$ atoms are in hyperfine state $A$ and $N_{B}$ atoms in hyperfine state $B$. The atoms interact via contact interactions of strengths $U_{ss^{\prime}}$ where $s$ and $s^{\prime}$ label the components $A$ and $B$. We assume that $N_{B}<N_{A}$, so that $B$ is by definition the minority component, and that the two components have the same mass $M$. Within a mean-field Gross-Pitaevskii (GP) description, the yrast spectrum (in units of $N\hbar^{2}/2MR^{2}$) can be calculated by
@@
Figure 2: The internal part of the mean-field yrast spectrum of an SU(2)-symmetric system, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=100$. The circles indicate the locations at which the derivative of the spectrum is discontinuous.
@@
whereas the yrast state is generally a soliton state at other fractional values of $l$. Consequently, the critical curves $x_{B}(\gamma,k)$ in Fig. 3 can also be interpreted as delineating the phases in which different numbers of plane-wave yrast states appear. It is interesting to see how these plane-wave yrast states emerge from soliton states as the interaction strength changes. For example, consider a system with $x_{B}=0.02$ and increase $\gamma$ from a small value (see the dashed line with an arrow in Fig. ). Since this value of $x_{B}$ lies within the range
@@
with $k=3$, the yrast spectrum has at most three derivative discontinuities in the fundamental range. As noted earlier, the yrast state at $l=x_{B}$ is always $(\phi_{0},\phi_{1})$, independent of $\gamma$. The yrast spectrum develops a second derivative discontinuity at $l=2x_{B}$ once $\gamma>9$, a value determined by the intersection of the line $x_{B}=0.02$ with the critical curve $x_{B}(\gamma,2)$. As $\gamma$ approaches this critical value from below, the soliton state at $l=2x_{B}$ continuously evolves into the plane-wave state $(\phi_{0},\phi_{2})$. When $\gamma$ reaches the critical value, this plane-wave state emerges and becomes pinned at $l=2x_{B}$ as $\gamma$ increases further. Finally, a third and final plane-wave yrast state $(\phi_{0},\phi_{3})$ emerges at $l=3x_{B}$ when $\gamma>60$, as determined by the intersection of the same line with the critical curve $x_{B}(\gamma,3)$.
The above summary shows that, for the SU(2)-symmetric system, the yrast spectrum in the fundamental range consists of smooth soliton branches joined by plane-wave states $(\phi_{0},\phi_{k})$ at fractional values $l=kx_{B}$, where each junction appears as a non-analytic point. Owning to the periodicity and the inversion symmetry of the internal energy, the appearance of the plane-wave yrast state $(\phi_{0},\phi_{k})$ necessarily implies the simultaneous appearance of the states $(\phi_{\mu},\phi_{\mu\pm k})$ where $\mu=1,2,\cdots$. In other words, the analytic structure of the yrast spectrum is fundamentally determined by the emergence of these plane-wave yrast states. We can therefore state our problem more concisely: for the asymmetric system, which plane-wave states can become the yrast states, and under what conditions? Or, equivalently, what are the critical curves and phase diagram of the plane-wave yrast states for the asymmetric system? In addressing this question using numerical solutions of the GP equations, we will also determine whether the emergence of plane-wave yrast states from soliton states always occurs in a continuous manner.
@@ III Numerical method
Figure 5: The internal part of the mean-field yrast spectrum of an asymmetric system with $\kappa=1$, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=20$. The circles indicate the locations at which the derivative of the spectrum is discontinuous.
@@
Figure 7: Critical curves for the emergence of plane-wave yrast states and the corresponding phase diagram for $\kappa=-0.1$. Here value of $K$ indicates the total number of plane-wave yrast states in that phase.
In Fig. 4, we present examples of the internal yrast spectrum calculated for several values of $x_{B}$ at $\kappa=-0.1$ and $\gamma=20$. We observe that the spectrum exhibits no derivative discontinuity for $x_{B}=0.1$, a single discontinuity at $l=0.05$ for $x_{B}=0.05$, and two discontinuities at $l=0.02$ and $l=0.04$ for $x_{B}=0.02$. Similar results for $\kappa=1$ and $\gamma=20$ are shown in Fig. 5, where the spectrum exhibits one derivative discontinuity at $l=0.4$ for $x_{B}=0.4$, two discontinuities at $l=0.2$ and $l=0.4$ for $x_{B}=0.2$, and three discontinuities at $l=0.1$, $0.2$, and $0.3$ for $x_{B}=0.1$. At all of these discontinuities, located at $l=kx_{B}$, the condensate wave functions correspond to the plane-wave states $(\phi_{0},\phi_{k})$. All of these features of the yrast spectrum will become transparent once the critical conditions for the emergence of plane-wave yrast states, discussed below, are established.
@@ IV Critical conditions and phase diagrams
Figure 8: The upper and middle panels, respectively, show the amplitude and phase of the condensate wave function at $l=2x_{B}$ with $x_{B}=0.02$ for several values of $\gamma$. The continuous transition from the soliton state to the plane-wave state $(\phi_{0},\phi_{2})$ is most clearly illustrated by a plot of the amplitude as a function of the phase in the polar coordinate, displayed in the bottom panel.
It is clear that the critical curve $x_{B}(\gamma,k)$ for the emergence of the $(\phi_{0},\phi_{k})$ yrast state depends on the parameter $\kappa$ characterizing the interaction asymmetry. Because of the inversion symmetry we can restrict the possible $k$ to positive integers. We now present a simple argument showing that the behavior of these critical curves differ qualitatively for $\kappa<0$ and $\kappa>0$. From Eq. (2) we find that the energy of the plane-wave state $(\phi_{0},\phi_{k})$ is given by
@@
with $\delta\rho_{s}(\theta)=\rho_{s}(\theta)-1/2\pi$ and $\delta\rho(\theta)=x_{A}\delta\rho_{A}(\theta)+x_{B}\delta\rho_{B}(\theta)$. Now suppose that the plane-wave state $(\phi_{0},\phi_{k})$ is an yrast state at $\kappa=0$ for some parameter set $(\gamma,x_{B})$, implying that it has the lowest energy in the $l=kx_{B}$ manifold, i.e.,
@@
The above analysis shows that an asymmetric interaction with $\kappa>0$ stabilizes the plane-wave state as an yrast state, whereas $\kappa<0$ tends to destabilize it. Given the saturating behavior of the critical curves at $\kappa=0$ shown in Fig. , we expect the curves for $\kappa<0$ to tilt downward and move away from the asymptotic lines, while those for $\kappa>0$ tilt upward and cross these lines.
@@ IV.1 $\kappa<0$
As an example, we show in Fig. 6 the critical curve for the plane-wave yrast state $(\phi_{0},\phi_{2})$ at $\kappa=-0.1$, alongside the corresponding curves for $\kappa=0$. The curve at $\kappa=-0.1$ behaves exactly as anticipated from the preceding argument. In particular, the critical values of $x_{B}$ all approach zero at large $\gamma$. Since the plane-wave state $(\phi_{0},\phi_{k})$ becomes an yrast state in the region bounded by its critical curve and the $x_{B}=0$ axis, this implies that the state will eventually cease to be an yrast state at sufficiently large $\gamma$, regardless of the value of $x_{B}$. Such a behavior can also be understood from Eqs. (21) and (22), which show that $\delta\bar{E}_{0}(\kappa<0)$ decreases without bound as $\gamma$ increases and will ultimately become negative. In other words, for a fixed negative value of $\kappa$, increasing $\gamma$ also tends to destabilize the plane-wave state as an yrast state. This is most clearly illustrated by the $(\phi_{0},\phi_{1})$ state: it is always an yrast state at $\kappa=0$, yet for $\kappa<0$ it ceases to be one once $\gamma$ exceeds a certain threshold (see Fig. 7). In fact, the entire system becomes unstable once $\gamma$ exceeds values set by the dynamical stability condition
@@
The stability curve together with a partial phase diagram of the plane-wave yrast states is shown in Fig. 7 for $\kappa=-0.1$. A complete phase diagram is not feasible to present, as phases hosting an increasing number of plane-wave yrast states emerge at progressively smaller values of $x_{B}$ and increasingly larger values of $\gamma$.
Interestingly, we find that the critical curves for $\kappa<0$ obtained from numerical solutions of the coupled GP equations agree with those determined using a perturbative approach. As discussed in Sec. I, this implies that the assumption of a continuous transition between the soliton states and the plane-wave yrast states at $l=kx_{B}$ is correct in this case. This is verified by our numerical calculations. As a representative example, we fix $x_{B}=0.02$ and increase $\gamma$. As shown in the phase diagram in Fig. 6, the plane-wave state $(\phi_{0},\phi_{2})$ becomes an yrast state once $\gamma$ exceeds the critical value $\gamma_{c}\simeq 11.2$. In Fig. 8, we plot the condensate wave functions at $l=2x_{B}$ for $\gamma=8$, $10$, $11$, and $11.5$. These results clearly demonstrate a continuous evolution from soliton states to the plane-wave state $(\phi_{0},\phi_{2})$ as $\gamma$ approaches and crosses the critical value.

### v2

@@ Abstract
The mean-field yrast spectrum of an SU(2)-symmetric two-component Bose gas confined to a ring geometry is known to exhibit an intricate nonanalytic structure that is absent in single-component systems. In particular, due to the interplay between the species concentration and the atomic interactions, a sequence of plane-wave states can emerge as yrast states at fractional values of the angular momentum per particle. This behavior stands in sharp contrast to the single-component case, where plane-wave states occur only at integer angular momenta. In this paper, we investigate how the structure of the yrast spectrum in a two-component Bose gas is modified by interaction asymmetry. By numerically solving the coupled Gross-Pitaevskii equations for propagating soliton states, we compute the mean-field yrast spectrum and, in particular, determine the critical curves associated with the emergence of various plane-wave yrast states. We find that both the behavior of these critical curves and the mechanisms by which plane-wave yrast states arise depend sensitively on the relative strengths of the inter- and intra-component interactions. When the inter-component interaction is weaker, the plane-wave yrast states replace soliton states through a continuous evolution, as in the SU(2)-symmetric case, although the conditions for their existence become more restrictive. In contrast, when the inter-component interaction is stronger, plane-wave yrast states emerge by overtaking soliton states via branch crossings, and their stability is significantly enhanced. Our results have important implications for the existence and stability of persistent currents in asymmetric, two-component Bose gases.
@@ I Introduction
In recent decades, quantum gases have become as an important platform for investigating persistent currents for two principal reasons [Ryu07, Ramanathan11, Moulder12, Beattie13, Wright13, Cai22, Del22, Polo25]. First, experimental advances in realizing multicomponent superfluids and synthetic gauge fields have substantially broadened the class of systems that may support persistent currents [KAWAGUCHI12, Lin_2016]. Second, the yrast spectrum of such systems can be determined quantitatively by mean-field theory, revealing a rich structure and enabling persistent currents beyond the conventional single-component scenario [Smyrnakis09, Anoshkin2013, Wu2013, Abad14, Wu2015, Pecci23]. Indeed, the mean-field yrast spectrum of a SU(2)-symmetric two-component Bose gas in a ring geometry is known to exhibit an intricate structure not present in single-component systems [Smyrnakis09, Anoshkin2013, Wu2013]. Although it still obeys the general form given in Eq. (1), the yrast spectrum of this two-component system can develop a series of non-analytic points at certain fractional values of $l$, arising from the interplay between component concentration and atomic interactions. The states at these additional points, much like those at integer $l$, may also support persistent currents. This feature of the yrast spectrum was first uncovered using analytic soliton solutions of the coupled Gross-Pitaevskii equations for the SU(2)-symmetric system, and its implications for persistent currents were later confirmed experimentally [Beattie13]. Notably, the appearance of each additional non-analytic point at the fractional $l$ is accompanied by a transition in the condensate wave functions: a pair of soliton solutions is replaced by a pair of plane-wave states. Moreover, these plane-wave states carry different winding numbers, reflecting distinct quantizations of circulation in the two components.
@@
Most experimentally accessible two-component systems are not SU(2) symmetric, since the intra- and inter-component interaction strengths generally differ. In this case, the coupled Gross-Pitaevskii (GP) equations no longer admit analytic solutions, and it is therefore natural to ask whether the recently discovered features in the yrast spectrum persist under these more realistic conditions. Reference [Wu2015] attempted to address this question using a perturbative approach based on an assumption motivated by the SU(2)-symmetric case, namely, that the appearance of a nonanalytic point at a fractional value of the angular momentum $l$ is always accompanied by a continuous transformation of the corresponding soliton state into a plane-wave state at that angular momentum. However, this assumption has never been verified for asymmetric systems. Several other works [Smyrnakis2014, Sandin16, Matsushita2018, Roussou2018] have also numerically investigated the yrast spectrum of asymmetric two-component systems, but an overarching picture of the structure of the yrast spectrum has yet to emerge. In this work, we present a comprehensive study of the mean-field yrast spectrum of asymmetric two-component Bose gases confined to a ring geometry by numerically solving the coupled GP equations. We find that the assumption adopted in Ref. [Wu2015] holds only for asymmetric two-component gases in which the inter-component interaction is weaker than the intra-component interaction. In the opposite regime, this assumption breaks down: the plane-wave yrast states instead emerge from crossings between distinct branches of solutions to the GP equations, and the resulting yrast spectrum exhibits an even richer nonanalytic structure than in the SU(2)-symmetric case. In both regimes, we establish critical conditions for the emergence of plane-wave yrast states, which to a large extent determine the analytic structure of the yrast spectrum.
@@ II Statement of the problem
We consider a two-component Bose gas of $N$ atoms confined to a ring of radius $R$, where $N_{A}$ atoms are in hyperfine state $A$ and $N_{B}$ atoms in hyperfine state $B$. The atoms interact via contact interactions of strengths $U_{ss^{\prime}}$ where $s$ and $s^{\prime}$ label the components $A$ and $B$. We assume that $N_{B}<N_{A}$, so that $B$ is by definition the minority component, and that the two components have the same mass $M$. Within a mean-field Gross-Pitaevskii description, the yrast spectrum (in units of $N\hbar^{2}/2MR^{2}$) can be calculated by
@@
Figure 2: The internal part of the mean-field yrast spectrum of an SU(2)-symmetric system, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=100$. The circles mark the points at which the derivative of the spectrum is discontinuous and at which the corresponding yrast state is a plane-wave state.
@@
whereas the yrast state is generally a soliton state at other fractional values of $l$. Consequently, the critical curves $x_{B}(\gamma,k)$ in Fig. 3 can also be interpreted as delineating the phases in which different numbers of plane-wave yrast states appear. It is interesting to see how these plane-wave yrast states emerge from soliton states as the interaction strength changes. For example, consider a system with $x_{B}=0.02$ and increase $\gamma$ from a small value (see the dashed line with an arrow in Fig. 3). Since this value of $x_{B}$ lies within the range
@@
with $k=3$, the yrast spectrum has at most three derivative discontinuities in the fundamental range. As noted earlier, the yrast state at $l=x_{B}$ is always $(\phi_{0},\phi_{1})$, independent of $\gamma$. The yrast spectrum develops a second derivative discontinuity at $l=2x_{B}$ once $\gamma\gtrsim 8.1$, a value determined by the intersection of the line $x_{B}=0.02$ with the critical curve $x_{B}(\gamma,2)$. As $\gamma$ approaches this critical value from below, the soliton state at $l=2x_{B}$ continuously evolves into the plane-wave state $(\phi_{0},\phi_{2})$. When $\gamma$ reaches the critical value, this plane-wave state emerges and becomes pinned at $l=2x_{B}$ as $\gamma$ increases further. Finally, a third and final plane-wave yrast state $(\phi_{0},\phi_{3})$ emerges at $l=3x_{B}$ when $\gamma\gtrsim 54.2$, as determined by the intersection of the same line with the critical curve $x_{B}(\gamma,3)$.
The above summary shows that, for the SU(2)-symmetric system, the yrast spectrum in the fundamental range consists of smooth soliton branches joined by plane-wave states $(\phi_{0},\phi_{k})$ at fractional values $l=kx_{B}$, where each junction appears as a nonanalytic point. Owing to the periodicity and the inversion symmetry of the internal energy, the appearance of the plane-wave yrast state $(\phi_{0},\phi_{k})$ necessarily implies the simultaneous appearance of the states $(\phi_{\mu},\phi_{\mu\pm k})$ where $\mu=1,2,\cdots$. In other words, the analytic structure of the yrast spectrum is fundamentally determined by the emergence of these plane-wave yrast states. We can therefore state our problem more concisely: for the asymmetric system, which plane-wave states can become the yrast states, and under what conditions? Or, equivalently, what are the critical curves and phase diagram of the plane-wave yrast states for the asymmetric system? In addressing this question using numerical solutions of the GP equations, we will also determine whether the emergence of plane-wave yrast states from soliton states always occurs in a continuous manner.
@@ III Numerical method
Figure 5: The internal part of the mean-field yrast spectrum of an asymmetric system with $\kappa=1$, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=20$. The circles mark the points at which the derivative of the spectrum is discontinuous and at which the corresponding yrast state is a plane-wave state.
@@
Figure 7: The critical curves for the emergence of plane-wave yrast states and the corresponding phase diagram for $\kappa=-0.1$. Here value of $K$ indicates the total number of plane-wave yrast states in that phase.
In Fig. 4, we present examples of the internal yrast spectrum calculated for several values of $x_{B}$ at $\kappa=-0.1$ and $\gamma=20$. We observe that the spectrum exhibits no derivative discontinuity for $x_{B}=0.1$, a single discontinuity at $l=0.05$ for $x_{B}=0.05$, and two discontinuities at $l=0.02$ and $l=0.04$ for $x_{B}=0.02$. Similar results for $\kappa=1$ and $\gamma=20$ are shown in Fig. 5, where the spectrum exhibits one derivative discontinuity at $l=0.4$ for $x_{B}=0.4$, two discontinuities at $l=0.2$ and $l=0.4$ for $x_{B}=0.2$, and three discontinuities at $l=0.15$, $0.3$, and $0.45$ for $x_{B}=0.15$. At all of these discontinuities, located at $l=kx_{B}$, the condensate wave functions correspond to the plane-wave states $(\phi_{0},\phi_{k})$. All of these features of the yrast spectrum will become transparent once the critical conditions for the emergence of plane-wave yrast states, discussed below, are established.
@@ IV Critical conditions and phase diagrams
Figure 8: The upper and middle panels, respectively, show the amplitude and phase of the condensate wave function at $l=2x_{B}$ with $x_{B}=0.02$ and $\kappa=-0.1$ for $\gamma=8,10,11$ and $11.5$ (see the dots on the dashed line with arrow in Fig. 7). The continuous transition from the soliton state to the plane-wave state $(\phi_{0},\phi_{2})$ as $\gamma$ crosses the critical curve of $k=2$ is most clearly illustrated by a plot of the amplitude as a function of the phase in the polar coordinate, displayed in the bottom panel.
To determine the critical curve $x_{B}(\gamma,k)$ for the emergence of the $(\phi_{0},\phi_{k})$ yrast state, we can restrict the possible $k$ to positive integers because of the inversion symmetry of the yrast spectrum. It is clear that these curves also depend on the parameter $\kappa$ characterizing the interaction asymmetry. We now present a simple argument showing that the behavior of these critical curves differ qualitatively for $\kappa<0$ and $\kappa>0$. From Eq. (2) we find that the energy of the plane-wave state $(\phi_{0},\phi_{k})$ is given by
@@
with $\delta\rho_{s}(\theta)=\rho_{s}(\theta)-1/2\pi$ and $\delta\rho(\theta)=x_{A}\delta\rho_{A}(\theta)+x_{B}\delta\rho_{B}(\theta)$. Now suppose that the plane-wave state $(\phi_{0},\phi_{k})$ is an yrast state at $\kappa=0$ for some parameter set $(\gamma,x_{B})$, meaning that it has the lowest energy in the $l=kx_{B}$ manifold, i.e.,
@@
The above analysis shows that an asymmetric interaction with $\kappa>0$ stabilizes the plane-wave state as an yrast state, whereas $\kappa<0$ tends to destabilize it. Given the saturating behavior of the critical curves at $\kappa=0$ shown in Fig. 3, we expect the curves for $\kappa<0$ to tilt downward and move away from the asymptotic lines, while those for $\kappa>0$ tilt upward and cross these lines.
@@ IV.1 $\kappa<0$
As an example, we show in Fig. 6 the critical curve for the plane-wave yrast state $(\phi_{0},\phi_{2})$ at $\kappa=-0.1$, alongside the corresponding curves for $\kappa=0$. The curve at $\kappa=-0.1$ behaves exactly as anticipated from the preceding argument. In particular, the critical values of $x_{B}$ approach zero at large $\gamma$. Since the plane-wave state $(\phi_{0},\phi_{k})$ becomes an yrast state in the region bounded by its critical curve and the $x_{B}=0$ axis, this implies that this state, after becoming an yrast state, will eventually cease to be an yrast state at sufficiently large $\gamma$, regardless of the value of $x_{B}$. Such a behavior can also be understood from Eqs. (21) and (22), which show that $\delta\bar{E}_{0}(\kappa<0)$ decreases without bound as $\gamma$ increases and will ultimately become negative. In other words, for a fixed negative value of $\kappa$, increasing $\gamma$ tends to destabilize the plane-wave state as an yrast state. This is most clearly illustrated by the $(\phi_{0},\phi_{1})$ state: it is always an yrast state at $\kappa=0$, yet for $\kappa<0$ it ceases to be one once $\gamma$ exceeds a certain threshold (see Fig. 7). In fact, the entire system becomes unstable once $\gamma$ exceeds values set by the dynamical stability condition
@@
The stability curve together with a partial phase diagram of the plane-wave yrast states is shown in Fig. 7 for $\kappa=-0.1$. A complete phase diagram is not feasible to present, as phases hosting an increasing number of plane-wave yrast states emerge at progressively smaller values of $x_{B}$ and increasingly larger values of $\gamma$. The nonanalytic structure of the yrast spectrum shown in Fig. 4 can now be understood in light of the corresponding phase diagram. Specifically, for a given set of parameters $(x_{B},\gamma)$, the number of derivative discontinuities in the spectrum associated with plane-wave yrast states can be predicated by the location of $(x_{B},\gamma)$ within the phase diagram. For instance, the yrast spectrum for $(x_{B},\gamma)=(0.02,20)$ has two derivative discontinuities at $l=x_{B}$ and $2x_{B}$ because the parameter set lies in the $K=2$ phase.
Interestingly, we find that the critical curves for $\kappa<0$ obtained from numerical solutions of the coupled GP equations agree with those determined by Ref. [Wu2015] using a perturbative approach. As discussed in Sec. I, this implies that the assumption of a continuous transition between the soliton states and the plane-wave yrast states at $l=kx_{B}$ is correct in this case. This is verified by our numerical calculations. As a representative example, we fix $x_{B}=0.02$ and increase $\gamma$. As shown in the phase diagram in Fig. 6, the plane-wave state $(\phi_{0},\phi_{2})$ becomes an yrast state once $\gamma$ exceeds the critical value $\gamma_{c}\simeq 11.2$. In Fig. 8, we plot the condensate wave functions at $l=2x_{B}$ for $\gamma=8$, $10$, $11$, and $11.5$. These results clearly demonstrate a continuous evolution from soliton states to the plane-wave state $(\phi_{0},\phi_{2})$ as $\gamma$ approaches and crosses the critical value.